OpenStudy (anonymous):

26. The perimeter of a rectangular concrete slab is 114 feet and its area is 702 square feet. Find the dimensions of the rectangle. a. Using l for the length of the rectangle, write an expression for the width of the rectangle in terms of l. (Hint: Solve the formula for w.) Show your work. b. Write a quadratic equation using l, the expression you found in part (a), and the area of the slab. c. Solve the quadratic equation. Use the two solutions to find the dimensions of the rectangle. 5 medals for CORRECT answer!!

OpenStudy (anonymous):

5 medals or more

OpenStudy (hoblos):

L+w = 114 so w=114-L Lw=702 L(114-l)=702 -L^2 + 114L =702 L^2 - 114L +702 do you know how to solve a quadratic?

OpenStudy (anonymous):

I'm not an expert. Is this correct? HONESTY IS THE BEST FOR A PERSON. -Aaron

OpenStudy (anonymous):

Isn't it this..

OpenStudy (anonymous):

PART C?

OpenStudy (hoblos):

i gave you parts a &b now we have to solve the quadratic equation just apply the formula do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please help dude!!

OpenStudy (hoblos):

actually i forgot it!! im sorry search for it you must have taken it in school

OpenStudy (anonymous):

darn it